# {textblock}

### v1

@@ Abstract
Background. The recent surge in phishing attacks keeps undermining the effectiveness of the traditional anti-phishing blacklist approaches. On-device anti-phishing solutions are gaining popularity as they offer faster phishing detection locally.
Aim. We aim to eliminate the delay in recognizing and recording phishing campaigns in databases via on-device solutions that identify phishing sites immediately when encountered by the user rather than waiting for a web crawler’s scan to finish. Additionally, utilizing operating system-specific resources and frameworks, we aim to minimize the impact on system performance and depend on local processing to protect user privacy.
Method. We propose a phishing detection solution that uses a combination of computer vision and on-device machine learning models to analyze websites in real time. Our reference-based approach analyzes the visual content of webpages, identifying phishing attempts through layout analysis, credential input areas detection, and brand impersonation criteria combination.
Results. Our case study shows it’s feasible to perform background processing on-device continuously, for the case of the web browser requiring the resource use of 16% of a single CPU core and less than 84MB of RAM on Apple M1 while maintaining the accuracy of brand logo detection at 46.6% (comparable with baselines), and of Credential Requiring Page detection at 98.1% (improving the baseline by 3.1%), within the test dataset.
Conclusions.  Our results demonstrate the potential of on-device, real-time phishing detection systems to enhance cybersecurity defensive technologies and extend the scope of phishing detection to more similar regions of interest, e.g., email clients and messenger windows.
@@ 1 Introduction
According to the phishing activity trends reports [^1] https://apwg.org/trendsreports/ from the Anti-Phishing Working Group (APWG), the number of phishing attacks keeps increasing year over year, with 2023 being the worst so far, with 4,987,809 unique phishing webpages created. While cloud infrastructure providers like Cloudflare implement phishing detection on their end [^2] https://blog.cloudflare.com/2023-phishing-report, these measures are not sufficient to prevent phishing attacks at large, so there is still a need for on-device solutions, both for desktop and mobile.
Existing phishing detection methods can be categorized into three categories: blacklist-based, classification-based, and reference-based. Traditional blacklist-based solutions like Google Safe Browsing  [^3] https://safebrowsing.google.com, although they are effective, do not keep up with the speed of phishing websites spreading [19], whose creation and deployment are typically automated. Classification-based approaches [18, 17, 10] use machine learning algorithms to analyze URL, HTML, or other features to classify webpages as phishing or legitimate, including on-device approaches [7, 2], but lose efficiency against HTML obfuscation techniques and lack accuracy. In contrast, reference-based approaches [1, 13, 14] use computer vision and can effectively analyze webpage appearance to extract information about the displayed content, allowing to detect phishing attempts and form reasonable verdicts.
@@ 2 Threat Model
Figure 1: General workflow of our solution
@@ 3 Solution
Figure 1 presents a high-level overview of our proposed approach. Similarly to PhishIntention, we focus on identifying both the intention to capture credentials and the intention to impersonate a brand. We begin with the webpage layout analysis (Section 3.1). In the next step, we classify the webpage elements into two categories: credential required layout or not (Section 3.2). The final step is to classify the brand that the phishing webpage is trying to impersonate well-known brand (Section 3.3).
Data collected in these steps will help us to verify all three conditions: (a) verifying that the webpage layout is designed to collect credentials, (b) confirming that a detected brand appears on our reference list, and (c) ensuring that the webpage does not belong to one of the official brand domains. Matching on all three conditions allows us to conclude that the current page is attempting to phish users.
@@ 3.1 Abstract Layout Detection
The main purpose of abstract layout detection is to determine whether the currently visible webpage contains credential input forms that may be targeted by hackers and find the brand logo bounding box. To achieve this, we analyze screenshots, identifying different layout elements such as logos, buttons, inputs, labels, and blocks.
We have used the part of the dataset from the section CRP classifier and AWL detector of [15]. The dataset contains about 9K webpage screenshots, annotated with regions and types of layout elements, together with the information about the webpage layout type (credential required or not) and the corresponding brand.
@@
The NMS (Non-Maximum Suppression) algorithm is used in object detection to eliminate redundant or overlapping bounding boxes. By design, the DETR model doesn’t have a built-in NMS layer. However, given that an abstract webpage layout should never contain overlapping elements, we integrated the protobuf model from coremltools  [^5] https://apple.github.io/coremltools/v3.4/index.html and merged it with the main model. DETR consistently generates a fixed set of 100 bounding boxes for any input image, which is the default value but can be changed. These bounding boxes are a representation of potential object locations within the image (Figure 2).
@@ 3.2 Credential Required Page Classification
When we have successfully extracted the layout from the webpage image, we also need to classify this layout into credential or non-credential required classes. To streamline our process and optimize algorithm we aimed to utilize already extracted features from Abstract Layout Detection step. The DETR model receives the query image as an input and generates N output vectors. Because output vectors are used to predict the class and location of the object, they consist of the information about the object and are suitable for the object-level image representation [4].
The object detection model is now divided into two models, the first being LayoutFeatureExtractor, which generates object-level features of shape (B, N, H) where B is batch size, N is a number of output vectors which is 100 in our case and H is hidden states for each output which is 256 in our case. The subsequent LayoutObjectDetector model takes these raw features as input and transforms it into a tensor of shape (B, N, 5) where the first 4 elements are coordinates of the bounding boxes and the last element is a layout element class. Then, we employ an additional LayoutClassifier model, which repurposes the previously extracted object-level features to transform the task of object detection into image classification. This transition is facilitated by a custom classification head (Figure 4), comprising a streamlined multi-layer perceptron architecture. The classification head processes and flattens the object-level features, applies ReLU activation functions for non-linear processing, and concludes with a Dropout layer to mitigate overfitting. Resulting in the output of shape (B, 2), this classification head efficiently provides classification scores, distinctly categorizing pages as credential required or not.
@@
With frozen weights of LayoutFeatureExtractor, we exclusively trained the classification head to transform the extracted features into classification scores. This approach allowed us to save disk storage and reduce RAM consumption by reusing already extracted features and has also greatly improved accuracy (Section 4.1.2).
3.3 Logo Brand Classification
We use brand logos available in the CRP classifier and AWL detector dataset part of [15]. From the available brands in the dataset, we select 15 brands known worldwide. We crop the logo elements from the entire screenshot and associate them with the corresponding brands. We used the ResNet-18 model, which has been trained on precisely cropped logo layout elements during the training stage. During the evaluation stage, we utilize the layout elements identified by the LayoutObjectDetector (DETR) and select the logo element with the highest confidence score. Subsequently, the LogoBrandClassifier (ResNet-18) is applied within the bounds of the detected logo’s bounding box. Accuracy metrics details are available in the Section 4.1.3.
3.3.1 Phishing detection
We used Netify [^6] https://www.netlify.com/ to get primary domains for our detectable brands. This list doesn’t contain all the brands that we’re detecting, and some had to be collected manually. Engaging in this process provided profound insights into the mechanisms through which phishing operations can be successfully executed. For example, DHL does not provide a publicly accessible set of its official domains. This lack of transparency is compounded by the existence of multiple domains, such as dhl.com, express.dhl, mydhli.com, dhlsameday.com, and dhlexpresscommerce.com, which could potentially confuse users and expose them to phishing risks.
@@ 3.4 Models Hierarchy
All models were converted to the Core ML format to optimize performance through the use of Metal Performance Shaders. Core ML  [^7] https://docs.developer.apple.com/documentation/coreml is a built-in high-level macOS native framework to integrate machine learning models into Apple platform applications. It is optimized to perform on-device operations by leveraging the CPU, GPU, and Neural Engine chip while minimizing its memory footprint and power consumption. Given the need for models to adhere to the proprietary format (.mlmodel) for use, Core ML accommodates conversion from third-party frameworks through Model Intermediate Language  [^8] https://apple.github.io/coremltools/docs-guides/source/model-intermediate-language.html (MIL). Metal  [^9] https://docs.developer.apple.com/documentation/metal/ framework allows applications to directly interact with a device’s GPU. Machine learning applications exploit Metal’s computational acceleration for both training and inference tasks. Metal Performance Shaders  [^10] https://docs.developer.apple.com/documentation/metalperformanceshaders/ is a suite of highly optimized compute and graphics shaders engineered for seamless and efficient integration within Metal applications. These shaders are data-parallel primitives that are fine-tuned to exploit the unique hardware capabilities of each GPU family, maximizing performance.
@@ 4 Evaluation
We perform a case study with the Safari web browser, exploring several use aspects to ensure that our system is optimized for both performance and efficiency:
Abstract Layout Detection Efficiency. How does our model’s logo-detecting efficiency compare to the previous solutions in terms of both accuracy and processing speed?
Credential Required Page Classification Efficiency. What is the effectiveness of our object-level features-based classification compared to previous work?
Logo Brand Classification Efficiency. How accurately can we classify the brand of detected logo elements as part of a phishing detection system?
System Resources Usage. Can we perform continuous processing in the background without a major impact on user workflow?
@@ 4.1.1 Abstract Layout Detection
One of the main goals of layout detection is to find logo elements, so we have checked how well DETR does this task. The mAP (Mean Average Precision) measurement is under the IoU (Intersection over Union) thresholds [0.5:0.95] and shows slightly worse accuracy compared to the PhishIntention model, but at the same time, achieves the same results as Phishpedia [13] model (Table II). We achieved worse results than competitors, but our solution is more efficient in terms of processing speed.
@@ 4.1.2 Credential Required Page Classification
To validate the accuracy and robustness of object-level features based classification, we conducted a comparative analysis against the PhishIntention baseline (Table III). The higher test accuracy of our solution (98.1%) underscores its effectiveness in accurately classifying webpages, reflecting its robustness against overfitting and its enhanced ability to generalize from the training data to real-world applications. This increment (+3.1%) in test accuracy is particularly noteworthy as it indicates a significant reduction in false positives and false negatives, which are critical in the context of cybersecurity solutions.
TABLE III: CRP Classification Accuracy Comparison
@@
4.1.3 Logo Brand Classification
Determining a phishing website’s brand affiliation significantly depends on the accurate recognition and extraction of logos. This crucial phase of brand logo identification is substantially influenced by the performance of the underlying object detection model, particularly in accurately locating the logo’s bounding box. Table IV presents the accuracy metrics achieved by employing the ResNet-18 model for logo brand classification. These metrics illustrate the model’s performance in terms of its F1 score, precision, and recall. The combination of DETR for object detection and ResNet-18 model for image classification demonstrates good performance with an F1 score of 0.95, indicating a high degree of accuracy in the balanced measurement of precision and recall.
TABLE IV: Logo Brand Classification Accuracy
Solution F1 Precision Recall
DETR + ResNet-18 0.95 0.92 0.95
@@ 4.2 System Resources Usage
Quantization [^13] https://huggingface.co/docs/optimum/concept_guides/quantization is a process of clamping the model weights. We quantized all models from a default 32-bit floating point to a smaller but less accurate 16-bit floating point. This optimization technique effectively halves the model’s size while maintaining a minimal impact on its accuracy. Using model quantization (Table VII) yields substantial enhancements in frame processing time, CPU utilization, and, most notably, RAM consumption. Furthermore, this process significantly reduces our app bundle size, decreasing from 254 MB to 136 MB.
@@ 5 Related Work
In the original PhishIntention algorithm, Faster R-CNN [20] model was chosen, but we encountered different errors during attempts to convert this architecture to Core ML format. We have initially added support of a few operators to MIL, but after deeper investigation, it became clear that Faster R-CNN isn’t convertible to Core ML format because it requires a number of custom layers that are not available in Metal and Core ML, so we decided to continue with DETR architecture.
@@ 6 Discussion
One of the primary benefits of on-device processing is its ability to provide the fastest response to previously unknown threats without waiting for database updates. Our proposed solution can continuously perform processing in the background without extensive system resource usage. However, different hardware provides different levels of performance due to specific optimizations. An interesting finding was that the Intel processor demonstrates worse results in terms of CPU usage when the frame rate is not limited (Table V), but it becomes twice as efficient when the frame limit is set to the one (Table VI). This finding suggests we need to look more into how hardware acceleration works and test more devices to understand this behavior better.
@@ 7 Limitations
We only used a small subset of well-known brands for our research, which means our solution doesn’t protect users from phishing attacks for unknown brands. Such an approach also fails when the company rebrands. For example, from Facebook to Meta or from Twitter to X. In case of the need for a well-known brand logos database update, machine learning models should be retrained and updated on the user device. A recent study [11] has also shown a possible attack on techniques that rely on brand logo recognition, for which we are not robust.

### v2

# Position Paper: Think Globally, React Locally — Bringing Real-time Reference-based Website Phishing Detection on macOS

@@ Abstract
Background. The recent surge in phishing attacks keeps undermining the effectiveness of the existing anti-phishing approaches. While state-of-the-art phishing detection solutions can effectively detect phishing across the web, they mainly focus on automated web crawling and updating blacklists
Aim. The time after a phishing website is published on the web and remains unprocessed by detection systems is critical. We aim to reduce this time gap via real-time phishing detection solution and perform continuous background processing without extra user interaction
Method. We propose a real-time on-device solution that identifies phishing sites immediately when encountered by the user. Our reference-based approach analyzes the visual content of webpages, identifying phishing attempts through layout analysis, brand impersonation recognition, and credential input areas detection
Results. Our case study shows that it’s feasible to perform background processing on-device continuously. For web browser phishing detection, the process utilizes 16% of a single CPU core and less than 84MB of RAM on an Apple M1 while maintaining a high efficiency with 95.7% precision and 87.7% recall, based on a test dataset of 50K phishing and benign webpages
Conclusions.  Our results demonstrate the potential of on-device, real-time phishing detection systems to enhance cybersecurity defensive technologies. We maintained the accuracy of state-of-the-art reference-based phishing detection solutions while bringing this functionality directly to the device
@@ 1 Introduction
According to the phishing activity trends reports [^1] https://apwg.org/trendsreports/ from the Anti-Phishing Working Group (APWG), the number of phishing attacks keeps increasing year over year, with 2023 being the worst so far, with 4,987,809 unique phishing webpages created. While cloud infrastructure providers like Cloudflare implement phishing detection on their end [^2] https://blog.cloudflare.com/2023-phishing-report, these measures are not sufficient to prevent phishing attacks at large. Existing phishing detection methods can be categorized into three categories: blacklist-based, classification-based, and reference-based. Traditional blacklist-based solutions like Google Safe Browsing  [^3] https://safebrowsing.google.com, although they are effective, do not keep up with the speed of phishing websites spreading [19], whose creation and deployment are typically automated. Classification-based approaches [18, 17, 10] use machine learning algorithms to analyze URL, HTML, or other features to classify webpages as phishing or legitimate, including on-device approaches [7, 2], but lose efficiency against HTML obfuscation techniques and lack accuracy. In contrast, reference-based approaches [1, 13, 14] use computer vision and can effectively analyze webpage appearance to extract information about the displayed content, allowing to detect phishing attempts and form reasonable verdicts.
@@ 2 Threat Model
Figure 1: General workflow of our phishing detection approach
@@ 3 Solution
Figure 1 presents a high-level overview of our phishing detection approach. Similarly to PhishIntention, we focus on identifying both the intention to impersonate a brand and the intention to capture user personal information. The main algorithm consists of three essential steps. First, we begin with the webpage screenshot analysis (Section 3.1) to identify layout elements and gain rich knowledge about webpage content. Second, we classify the detected logo brand (Section 3.2) that the phishing webpage tries to impersonate. If we find a brand from our reference list, we continue to the next step. Third, we classify the webpage into two categories: if it requires credentials or not (Section 3.3). This step confirms that an impersonated phishing website is trying to steal personal user information. Matching these conditions allows us to decide whether a webpage is phishing or legitimate.
@@ 3.1 Abstract Layout Detection
The main purpose of abstract layout detection is to determine whether the currently visible webpage contains a brand logo that may be targeted by hackers and find credential input forms. To achieve this, we analyze screenshots, identifying different layout elements such as logos, buttons, inputs, labels, and blocks. We used the part of the dataset from the section CRP classifier and AWL detector of [15]. The dataset contains about 9K webpage screenshots, annotated with regions and types of layout elements, together with the information about the webpage layout type (credential required or not) and the corresponding brand.
@@
The NMS (Non-Maximum Suppression) algorithm is used in object detection to eliminate redundant or overlapping bounding boxes. By design, the DETR model doesn’t have a built-in NMS layer. However, given that an abstract webpage layout should never contain overlapping elements, we integrated the protobuf model from coremltools  [^5] https://apple.github.io/coremltools/docs-guides and merged it with the main model. DETR consistently generates a fixed set of 100 bounding boxes for any input image, which is the default value but can be changed. These bounding boxes are a representation of potential object locations within the image (Figure 2).
@@
3.2 Logo Brand Classification
To train logo brand classification model we used brand logos available in the OCR-aided Siamese model dataset part of [15]. This part contains 3061 logo images of 277 brands. We utilized the ResNet-18 model, which was trained on precisely cropped logo layout elements during the training stage. During the evaluation stage, we utilize the layout elements identified by the LayoutObjectDetector (DETR) and select the logo element with the highest confidence score. Subsequently, the LogoBrandClassifier (ResNet-18) is applied within the bounds of the detected logo’s bounding box. Accuracy metrics details are available in the Section 4.1.2.
@@ 3.3 Credential Required Page Classification
Credential required page classification is an important step to reduce the number of false positive detections. To streamline our process and optimize the algorithm we aimed to utilize already extracted features from the Abstract Layout Detection step. The DETR model receives the query image as an input and generates N output vectors. Because output vectors are used to predict the class and location of the object, they consist of the information about the object and are suitable for the object-level image representation [4]. The object detection model was divided into two models, the first is LayoutFeatureExtractor, which generates object-level features of shape (B, N, H) where B is batch size, N is a number of output vectors which is 100 in our case and H is hidden states for each output which is 256 in our case. The subsequent LayoutObjectDetector model takes these raw features as input and transforms it into a tensor of shape (B, N, 5) where the first 4 elements are coordinates of the bounding boxes and the last element is a layout element class. Then, we employ an additional LayoutClassifier model, which repurposes the previously extracted object-level features to transform the task of object detection into image classification. This transition is facilitated by a custom classification head (Figure 4), comprising a streamlined multi-layer perceptron architecture. The classification head processes and flattens the object-level features, applies ReLU activation functions for non-linear processing, and concludes with a Dropout layer to mitigate overfitting. Resulting in the output of shape (B, 2), this classification head efficiently provides classification scores, distinctly categorizing pages as credential required or not. Accuracy metrics details about this approach are available in the Section 4.1.3.
@@ 3.4 Models Hierarchy
All models were converted to the Core ML format to optimize performance through the use of Metal Performance Shaders. Core ML  [^6] https://docs.developer.apple.com/documentation/coreml is a built-in high-level macOS native framework to integrate machine learning models into Apple platform applications. It is optimized to perform on-device operations by leveraging the CPU, GPU, and Neural Engine chip while minimizing its memory footprint and power consumption. Metal  [^7] https://docs.developer.apple.com/documentation/metal/ framework allows applications to directly interact with a device’s GPU. Machine learning applications leverage Metal’s computational acceleration for both training and inference tasks.
@@ 4 Evaluation
We perform a case study, exploring several use aspects to ensure that our system is optimized for both performance and efficiency:
Abstract Layout Detection Accuracy. How does our model’s logo-detecting efficiency compare to the previous solutions in terms of both accuracy and processing speed?
Logo Brand Classification Accuracy. How accurately can we classify the brand of detected logo element?
Credential Required Page Classification Accuracy. What is the effectiveness of our object-level features-based classification compared to previous work?
Overall Phishing Detection Performance. Did we manage to maintain the overall accuracy of phishing webpage detection compared to other solutions?
System Resources Usage Efficiency. Can we perform continuous processing in the background without a major impact on user workflow?
@@ 4.1 Accuracy Metrics
During the accuracy evaluation, we used the same dataset [15] provided by PhishIntention team to compare our solution side by side with competitors. We provide detailed information about each used dataset part at the beginning of corresponding subsection.
@@ 4.1.1 Abstract Layout Detection
We used 901 labeled webpages from AWL detector dataset part of [15] to measure how accurate we can find logo elements. The mAP (Mean Average Precision) measurement is under the IoU (Intersection over Union) thresholds [0.5:0.95] and shows slightly worse accuracy compared to the PhishIntention model, but at the same time, achieves the same results as Phishpedia [13] model (Table II). We achieved worse results than competitors, but our solution is more efficient in terms of processing speed.
@@
4.1.2 Logo Brand Classification
We used 2000 logo images from OCR-aided Siamese model dataset part of  [15] to measure how accurate we can classify logo brands. Table III presents the accuracy metrics achieved by employing the ResNet-18 model for logo brand classification. We achieved the same level of accuracy as PhishIntention and slightly improved accuracy compared to Phishpedia.
TABLE III: Logo Brand Classification Accuracy
Solution Test Accuracy
PhishIntention 89.1
Phishpedia 83.5
Ours 90.8
@@ 4.1.3 Credential Required Page Classification
We used 901 labeled webpages from CRP classifier dataset part of [15] to validate the accuracy and robustness of object-level features based classification (Table IV). The increment (+3.1%) in test accuracy is particularly noteworthy as it indicates a reduction in false positive detections, which is clearly observed in Figure 7.
TABLE IV: CRP Classification Accuracy Comparison
@@
4.1.4 Overall Phishing Detection
Figure 6 shows the ROC (Receiver operating characteristic) of different phishing detection solutions in logarithm scale. We evaluated the algorithm within Experiment dataset (25K benign and 25K phishing webpages) part of [15]. This plot shows that our solution maintained phishing detection efficiency while bringing it directly to the device.
Figure 6: ROC of different phishing detection solutions
Figure 7 shows precision and recall within the same dataset, and additionally false positive rate within Misleading legitimacy dataset part of [15] which contains 3049 benign webpages. In this plot we can observe that increased accuracy of credential required page classification reduced the false positive rate from 5.1% to 3.4%.
Figure 7: Accuracy metrics of different phishing detection solutions
@@ 4.2 System Resources Usage
Quantization [^10] https://huggingface.co/docs/optimum/concept_guides/quantization is a process of clamping the model weights. We quantized all models from a default 32-bit floating point to a smaller but less accurate 16-bit floating point. Model quantization yields substantial enhancements in frame processing time, CPU utilization, bundle size, and, most notably, RAM consumption (Table VII).
@@
Subsequently, we checked how quantization affects overall phishing detection efficiency in the same way as described in Section 4.1.4. We found that this optimization technique effectively halves the model’s size while maintaining a minimal impact on its accuracy (Table VIII).
TABLE VIII: Quantization Impact on Accurracy
Floating-point Precision Recall
32-bit 95.8% 87.8%
16-bit 95.1% 87.1%
@@ 7 Limitations
Our solution doesn’t protect users from phishing attacks for unknown brands. In case of the need for a well-known brand logos database update, machine learning models should be retrained and updated on the user device. Such an approach also fails when the company rebrands. For example, from Facebook to Meta or from Twitter to X. A recent study [11] has also shown a possible attack on techniques that rely on brand logo recognition, for which we are not robust.